OpenStudy (anonymous):

please help. classify the system as independant, dependant, or inconsistant. problem in comments. best answer and i will become a fan.

OpenStudy (anonymous):

\[2y=x-4\]\[3x-6y=12\]

OpenStudy (anonymous):

lets multiply 2y=x-4 by 3 we get 6y=3x-12 which is the same as 3x-6y=12 so we have 3x-6y=0, so the system is dependent, that is for each x, an y exists so that the pair (x,y) satisfies your system.

OpenStudy (anonymous):

"so we have 3x-6y=0", was wrong line - it's 3x-6y=12

OpenStudy (anonymous):

thank you for the help lol even though im already a fan :)

OpenStudy (anonymous):

well, i am skilled in much more than linear systems, so don't be afraid to ask :P